OpenStudy (anonymous):

any one can help me to solve laplace tranform in RLC circuit?? Q(s) = 20/ s{(s+80)^2) + 3600} find q(t)

OpenStudy (anonymous):

is it... \[Q(s) = \frac{ 20 }{ s \left[ (s+80)^{2}+3600 \right] }\]

OpenStudy (anonymous):

That is what Mathematica gave \[ \mathcal{L}_s^{-1}\left[\frac{20}{s \left((s+80)^2+3600\right)}\right](x)=\\ 20 \left(\frac{1}{10000}+\frac{i e^{(-80-60 i) x} \left((4+3 i) e^{120 i x}+(-4+3 i)\right)}{60000}\right) \]